OpenStudy (anonymous):

DIVIDE 23x^5 u^3 +9x^4 u -12x^6 u^6/-3x^5 u^3

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (anonymous):

You have 3 expressions in the numerator, separated by a + or - sign The concept/idea here is to break up the fraction. Each term in the nukmerator gets divided by the denominator, so you get: 23x^5u^3/-3x^5u^3 + 9x^4u/-3x^5u^3 - 12x^6u^6/-3x^5u^3 Now we simplify each of the 3 fractions using law of exponents. Can you finish simplifying each fraction?

OpenStudy (anonymous):

23x^5u^3/-3x^5u^3

OpenStudy (anonymous):

is that the fraction

OpenStudy (anonymous):

Thats the first fraction....... That's equal to -23/3 (because the x^5 and u^3 cancel top and bottom) Now proceed to simplify the second fraction I wrote

OpenStudy (anonymous):

-1/3 (23 u^6 x^10)

OpenStudy (anonymous):

-3 u^4 x^9

OpenStudy (anonymous):

lets stop all this

OpenStudy (anonymous):

session is definitley timed out now

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

The second fraction is 9x^4u/-3x^5u^3 Simplify the numbers, then the x'sa, then the u's 9/-3 = -3 x^4/x^5..leaves you one x in the denominator u/u^3 leaves you a u^2 in the denominator so you get -3/xu^2

OpenStudy (anonymous):

Thats the second fraction. Now you simplify the third fraction.

OpenStudy (anonymous):

-4 u^9 x^11

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

Watch...the third fraction is: -12x^6u^6/-3x^5u^3 -12/-3 = +4 x^6/x^5 = x u^6/u^3 = u^3 so you get +4xu^3 Study what I just showed you and you tell me what you dont understand.

OpenStudy (anonymous):

whats the next step after simplifying the 3 fractions

OpenStudy (anonymous):

Put all three answers together You get: -23/3 - 3/xu^2 +4xu^3

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

welcome.